An investor pays interest at 10% p.a but the interest is paid quarterly. What is the effective annual rate? A 10.38% B

12.55% C 15.27% D 5%
Answer #1

Effective annual rate is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.10/4)^4 - 1

= 1.1038 - 1

= 0.1038*100

= 10.38%.

Hence, the answer is option a.

In case of any query, kindly comment on the solution.